# Uncovering Latent Themes of Messaging on Social Media by Integrating LLMs: A Case Study on Climate Campaigns

### v1

@@ Abstract
This paper introduces a novel approach to uncovering and analyzing themes in social media messaging. Recognizing the limitations of traditional topic-level analysis, which tends to capture only the overarching patterns, this study emphasizes the need for a finer-grained, theme-focused exploration. Conventional methods of theme discovery, involving manual processes and a human-in-the-loop approach, are valuable but face challenges in scalability, consistency, and resource intensity in terms of time and cost. To address these challenges, we propose a machine-in-the-loop approach that leverages the advanced capabilities of Large Language Models (LLMs). This approach allows for a deeper investigation into the thematic aspects of social media discourse, enabling us to uncover a diverse array of themes, each with unique characteristics and relevance, thereby offering a comprehensive understanding of the nuances present within broader topics. Furthermore, this method efficiently maps the text and the newly discovered themes, enhancing our understanding of the thematic nuances in social media messaging. We employ climate campaigns as a case study and demonstrate that our methodology yields more accurate and interpretable results compared to traditional topic models. Our results not only demonstrate the effectiveness of our approach in uncovering latent themes but also illuminate how these themes are tailored for demographic targeting in social media contexts. Additionally, our work sheds light on the dynamic nature of social media, revealing the shifts in the thematic focus of messaging in response to real-world events.
@@ Introduction
In recent years, social media has emerged as a fertile ground for diverse discussions and expressions, encompassing a broad spectrum of topics and themes. In this context, a ‘topic’ within social media content refers to a broad category under which various finer-grained ‘themes’ can be identified. The topic serves as an umbrella term that encompasses a range of more specific and nuanced themes. For example, under the ‘Climate Change’ topic, a theme might be ‘Renewable Energy Solutions’. Identifying themes within a topic involves delving into these specific aspects, revealing the detailed and granular nuances within the broader category. Given the dynamic nature of social media, accurately discerning and understanding the diverse themes within a broader topic is a complex and continuously evolving task.
@@
Traditionally, thematic analysis (TA) has been the method of choice for this task, serving as a qualitative research technique focused on pattern identification, where themes emerging from the data guide further analysis (Roberts, Dowell, and Nie 2019; Vaismoradi et al. 2016; Vaismoradi, Turunen, and Bondas 2013; Braun and Clarke 2012, 2006; Tuckett 2005). However, with the explosion of textual data in the digital age, there’s a growing inclination towards computational methods for text analysis (Sivarajah et al. 2017; Jagadish et al. 2014; Fan, Han, and Liu 2014).
@@
Among these, supervised text analysis methods like Support Vector Machines (SVMs) (Joachims 1998), Naive Bayes classifiers (McCallum, Nigam et al. 1998), and Neural Networks (LeCun, Bengio, and Hinton 2015) have been effective in theme categorization. These methods use pre-labeled datasets to learn and identify patterns, aiding in the categorization of new data. However, they present challenges, notably the need for extensive labeled datasets, which can be resource-intensive to create. Moreover, their performance heavily relies on the quality and representativeness of the training data. If the training data is biased or lacks diversity, the model’s ability to generalize and accurately identify themes in varied or new contexts is significantly compromised.
@@
On the other hand, unsupervised text analysis methods, such as Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003) and non-negative matrix factorization (NMF) (Lee and Seung 1999), have gained popularity in theme discovery. These methods independently identify themes from data without the need for pre-labeled datasets. While useful, they also have limitations. Topic modeling techniques like LDA can produce topics that are challenging to interpret or lack coherence, complicating the extraction of meaningful themes (Röder, Both, and Hinneburg 2015; Lau, Newman, and Baldwin 2014; Mimno et al. 2011; Chang et al. 2009). Recent advancements have seen the integration of human judgment through a human-in-the-loop approach, which, while enhancing accuracy, suffers from issues of scalability and high resource demand (Pacheco et al. 2022a, b; Lund et al. 2017; Hu et al. 2014). Recognizing this, our work explores a new approach, integrating the strengths of LLMs (Brown et al. 2020) with a streamlined algorithmic process to efficiently discover and characterize themes in social media messaging. This exploration leads us to the following research questions (RQ), which are crucial for assessing the potential of LLMs to enhance our analysis by providing deeper and more nuanced insights:
@@
To enhance theme discovery, we propose an iterative three-step machine-in-the-loop framework. We define large textual collections as repositories of textual instances (e.g., ads, tweets, posts, documents) where each instance is potentially associated with a set of annotated themes (seed set from Islam, Zhang, and Goldwasser (2023) that is used in this paper showed in Table 1). A high-level overview of our method is shown in Figure 1.
@@
To showcase our approach, we use climate campaigns (dataset released by Islam, Zhang, and Goldwasser (2023)) on social media as a case study. The dataset focuses on climate related English ads based on the United States from January $2021$ January $2022$ (Islam, Zhang, and Goldwasser 2023). Qualitative researchers use different theories to ground theme discovery, each associated with a different set of concepts. In this case, the theme discovery process is grounded using energy industry and climate change related stances (Islam, Zhang, and Goldwasser 2023). Our contributions are summarized as follows:
@@
    We conduct quantitative and qualitative analysis on real-world dataset to demonstrate the effectiveness of our proposed method. We show that our model outperforms the baselines.
@@
    Based on the newly discovered themes, we analyze demographic targeting on social media messaging.
@@
Pro-energy
Economy_pro, ClimateSolution, Pragmatism, Patriotism, AgainstClimatePolicy.
Clean-energy
Economy_clean, FutureGeneration, Environmental, HumanHealth, Animals, SupportClimatePolicy, AltEnergy, PoliticalAffiliation.
Table 1: Pre-existing themes from Islam, Zhang, and Goldwasser (2023).
@@
Several works have done for understanding messaging on social media (Islam, Roy, and Goldwasser 2023; Islam, Zhang, and Goldwasser 2023; Islam and Goldwasser 2022; Capozzi et al. 2021; Silva and Benevenuto 2021; Silva et al. 2020; Serrano et al. 2020; Mejova and Kalimeri 2020; Capozzi et al. 2020; Andreou et al. 2019). Most of the approaches to understanding messaging have relied on broader categories. However, in this paper, we focus on gaining a deeper understanding of messaging.
Previous works (Islam, Roy, and Goldwasser 2023; Islam, Zhang, and Goldwasser 2023; Islam and Goldwasser 2022; Capozzi et al. 2021) used a predefined set of labels, themes, and arguments to analyze messaging. These were fixed and established based on existing topics or theoretical frameworks, such as Moral Foundations Theory (MFT) (Haidt and Graham 2007). However, previous works often fall short of capturing the nuances of messaging choices. For instance, within the Economy topic, it may neglect to recognize conflicting arguments, such as Argument 1: “Increasing the minimum wage boosts economic growth,” versus Argument 2: “Increasing the minimum wage harms small businesses and reduces employment opportunities.” Recently, Pacheco et al. (2022a, b) suggested a human-in-the-loop approach to uncover new themes. In contrast to their work, we use LLMs inference.
Researchers in the fields of qualitative research (QR) and Natural Language Processing (NLP) are currently investigating how LLMs can be utilized for TA. Dai, Xiong, and Ku (2023) proposed LLM-in-the-loop approach for TA. Xiao et al. (2023) combined GPT-3 with expert-drafted codebooks for supporting qualitative analysis. Gao et al. (2023) suggested a collaborative interface for qualitative analysis in their work, utilizing LLMs for generating initial codes and aiding in decision-making processes. In this paper, we develop a new set of themes (or codes) specifically tailored for analyzing messaging in a particular domain, which helps speed up the development of domain-specific labels.
Framework
@@
Figure 2: Example of incoherent cluster.
@@
Figure 3: Example of merged cluster.
@@
Figure 4: Prompt template for coherency check (shown as zero-shot)
Figure 5: Prompt template for generating summary (shown as zero-shot))
Figure 6: Prompt template for LLMs Mapper (shown as four-shot)
@@
The second phase introduces an optional but crucial step – human validation. Two NLP and Computational Social Science (CSS) researchers did one $2$ hours session to assess the clusters formed by the algorithms. Their expertise is particularly invaluable in evaluating the nuances of language and context that automated systems might overlook. They scrutinize the merging decisions, the descriptiveness of the clusters, and overall coherence, bridging the gap between computational efficiency and the nuanced understanding of human judgment.
@@
In the final phase, we provide summaries of all clusters (including pre-existing + newly discovered) and prompt LLMs in a four-shot setting to assess whether a new instance aligns with the summarized clusters. To alleviate the issues of being hallucinated, we force LLMs to answer the questions by selecting from the given summarized clusters. The LLMs, equipped with the context from the initial phases, effectively categorize new texts into the relevant clusters (answering RQ2). This phase is crucial for maintaining the dynamism of the clusters, allowing the system to adapt to new data and emerging themes.
Method Num. Num. Covered Ads
Themes thr $<0.6$ thr $<0.5$ thr $<0.4$ thr $<0.3$
Pre-existing 13 14652 9725 3731 558
+After Iter1 20 18702 14583 8646 2944
+After Iter2 25 18988 15052 9079 3180
Table 2: Coverage. THR: Threshold, NUM: Number.
(a) Assignment with LLMs mapper.
(b) Assignment with SBERT.
Figure 7: Confusion Matrix after $2^{nd}$ round of iteration + assignment.
@@
(c) Ours: After $2^{nd}$ round of iteration.
Figure 8: Correlations between themes and stances.
Experiments
We use $21372$ ads and $13$ themes (Table 1) related to climate campaigns released by Islam, Zhang, and Goldwasser (2023). At first, we assign pre-existing themes to the ads using SBERT. Pre-existing themes can cover a small portion of ads ($1^{st}$ row of Table 2). The threshold is determined by the distance from the centroid to an ad, with a shorter distance being preferable. For a threshold of $<0.5$, there are $11647$ ads that remained uncovered, which we consider as a noisy set. We then apply k-means clustering (with $k=29$) to this noisy set. Following the candidate generation phase, we identify $7$ new themes in the $1^{st}$ iteration. We use GPT-3.5 for cluster coherency check, generating cluster summaries, and redundancy check. In $2^{nd}$ iteration, we repeat this process using $6789$ ads remaining from iteration 1 (again applying a threshold of $<0.5$), resulting in the discovery of $5$ additional themes. The complete set of final themes is presented in Table 3. Summaries of $25$ themes (clusters) are provided in Appendix A.
Iter1
BidenGasPriceIncrease, AgainstCorporateInterests, GasTax, Deforestation, Carbon, CustomerBasedAltEnergy, FoodSecurity
Iter2
EnergyAffordabilityandSustainabilityLegislation, EcofriendlyConsumerChoices, PlasticWasteandEnvironmentalImpact, PromoteSustainableTransportation, WaterManagementandSustainability
Table 3: Resulting Themes. Iter: Iteration.
Method Acc. Macro avg. F1
SBERT Assignment 84.05 79.32
LLMs Mapper (Ours) 88.15 89.24
Table 4: Mapping Quality w.r.t Human Judgements
Prompt Templates
This section shows the prompt templates used to check cluster coherency (Fig. 4), generate cluster summary (Fig. 5), and map ads to themes (Fig. 6).
Age Group
State
Entity
Claim
Theme
13-24 CA
Young People, Climate Change Denier, Climate Movement.
Importance of action against climate change to clean air & water and to reduce carbon emissions.
Environmental
TX
Natural Gas Systems, Coal Energy Systems, Nuclear Energy Systems, Wind Turbines, Solar Panels.
Failures in natural gas, coal, and nuclear systems caused almost double outages in Texas compared to frozen wind turbines and solar panels.
AltEnergy
25-54 CA
Carbon Emissions, Sacramento Municipal Utility District.
Eliminating all carbon emissions from the power supply by 2030.
Carbon
TX
Climate Change, Renewable Energy Plan.
Climate change impacts are irreversible by 2030, urging an immediate switch to renewable energy.
AltEnergy
55+ CA
Western Burrowing Owls, Aramis Industrial Solar Power Plant.
Aramis solar plant’s construction will harm wildlife and habitats, especially affecting Western burrowing owls.
Animals
TX
U.S. Senators or Congressman, Citizens’ Climate Lobby.
Urgent need for action on climate change.
SupClimPolicy
Table 5: Most mentioned entities, claim, and theme of targeted ads for young, working-age, older population of CA and TX. SupClimPolicy: SupportClimatePolicy.
Results and Analysis
Our experiments show that our framework can be used to uncover a set of themes that cover a large portion of the messaging regarding climate campaigns (Table 2). With a threshold $<0.4$, pre-existing themes can cover only $17.5\%$ ads. After $1^{st}$ iteration using our machine-in-loop approach, we cover $40.5\%$ and after $2^{nd}$ iteration, we cover $42.5\%$ ads (Table 2).
Gender
Entity
Claim
Theme
Male
Alex Askew, Karen Greenhalgh
Alex Askew’s record of raising taxes contrasted with Karen Greenhalgh’s promise to cut taxes and lower living costs.
GasTax
Female
Parents, Children/Kids
Urgent need for action against climate change to protect the future of children and the planet.
FutureGen
Table 6: Mentioned entities, claim, and theme of targeted ads for only male and only female population. FutureGen: FutureGeneration.
State Stance
Entity
Claim
Theme
VT Clean-energy Community
Climate change impacts health, family, and community.
HumanHealth
Pro-energy
Climate Action Council
Climate Action Council plans to increase fuel and carbon taxes.
GasTax
ND Pro-energy Community Oil and gas supports jobs and community growth. Economy_pro
Biden
Biden’s energy policies increase energy costs.
BidenGasPrInc
Clean-energy
Water system
Solution of water scarcity problem.
WaterMng&Sustain
Table 7: Entity, claim, and theme of targeted ads for blue and red states. The entity ‘Community’ is targeted differently by advertisers. While targeting the blue state Vermont, the emphasis is on the impacts of climate change on community health. In contrast, when targeting the red state North Dakota, the focus shifts to how oil and gas supports community growth. BidenGasPrInc: BidenGasPriceIncrease, WaterMng&Sustain: WaterManagementandSustainability.
@@
To evaluate the mapping quality of our model, we select ads that fall under threshold $<0.4$ and then randomly sample $1072$ ads. We ask the annotators to provide a score of $1$ if the ad aligns with the theme, 00 for no alignment. There is no clear definition of alignment, and we let the annotators make this judgment. To evaluate mapping quality, we compare the performance of LLMs mapper with SBERT assignment (Baseline) of ads to themes. Results are outlined in Table 4. The resulting mapping from ads to the theme is fairly accurate with respect to human judgments (Fig. 6(a)) with the accuracy of $88.15\%$ and macro avg. F1 score of $89.24\%$. Fig. 7 shows the confusion matrix after $2^{nd}$ round of iteration and assignment with LLMs mapper (6(a)) and SBERT mapper (6(b)).

### v2

# Discovering Latent Themes in Social Media Messaging: A Machine-in-the-Loop Approach Integrating LLMs

@@ Abstract
Grasping the themes of social media content is key to understanding the narratives that influence public opinion and behavior. The thematic analysis goes beyond traditional topic-level analysis, which often captures only the broadest patterns, providing deeper insights into specific and actionable themes such as “public sentiment towards vaccination”, “political discourse surrounding climate policies,” etc. In this paper, we introduce a novel approach to uncovering latent themes in social media messaging. Recognizing the limitations of the traditional topic-level analysis, which tends to capture only overarching patterns, this study emphasizes the need for a finer-grained, theme-focused exploration. Traditional theme discovery methods typically involve manual processes and a human-in-the-loop approach. While valuable, these methods face challenges in scalability, consistency, and resource intensity in terms of time and cost. To address these challenges, we propose a machine-in-the-loop approach that leverages the advanced capabilities of Large Language Models (LLMs). This approach facilitates a deeper investigation into the social media discourse, revealing a variety of themes with distinct characteristics and relevance. It provides a detailed understanding of underlying nuances and efficiently maps texts to these themes, enhancing our insight into social media messaging.
To demonstrate our approach, we apply our framework to contentious topics, such as climate debate and vaccine debate. We use two publicly available datasets: (1) the climate campaigns dataset of $21k$ Facebook ads and (2) the COVID-19 vaccine campaigns dataset of $9k$ Facebook ads. Our quantitative and qualitative analysis shows that our methodology yields more accurate and interpretable results compared to the baselines. Our results not only demonstrate the effectiveness of our approach in uncovering latent themes but also illuminate how these themes are tailored for demographic targeting in social media contexts. Additionally, our work sheds light on the dynamic nature of social media, revealing the shifts in the thematic focus of messaging in response to real-world events.
@@ Introduction
In recent years, social media has emerged as a fertile ground for diverse discussions and expressions, encompassing a broad spectrum of topics and themes. In this context, a ‘topic’ in social media content represents a broad category that includes various specific and nuanced ‘themes’. The topic acts as an umbrella term under which different themes fall. For example, under the ‘Climate Change’ topic, a theme might be ‘Renewable Energy Solutions’. Identifying themes within a topic is to explore this specific area in depth, thereby uncovering the detailed and granular nuances that define the broader category. Given the dynamic nature of social media, accurately discerning and understanding the diverse themes within a broader topic is a complex and continuously evolving task.
@@
Traditionally, thematic analysis (TA) has been the method of choice for this task, serving as a qualitative research technique focused on pattern identification, where themes emerging from the data guide further analysis (Roberts et al. 2019; Vaismoradi et al. 2016, 2013; Braun and Clarke 2012; Braun and et. al 2006; Tuckett 2005). However, with the explosion of textual data in the digital age, there’s a growing inclination towards computational methods for text analysis (Sivarajah et al. 2017; Jagadish et al. 2014; Fan et al. 2014).
@@
Among these, supervised text analysis methods like Support Vector Machines (SVMs) (Joachims 1998), Naive Bayes classifiers (McCallum et al. 1998), and Neural Networks (LeCun, Bengio, and Hinton 2015) have been effective in theme categorization. These methods use pre-labeled datasets to learn and identify patterns, aiding in the categorization of new data. However, they present challenges, notably the need for extensive labeled datasets, which can be resource-intensive to create. Moreover, their performance heavily relies on the quality and representativeness of the training data. If the training data is biased or lacks diversity, the model’s ability to generalize and accurately identify themes in varied or new contexts is significantly compromised.
@@
On the other hand, unsupervised text analysis methods, such as Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003) and non-negative matrix factorization (NMF) (Lee and Seung 1999), have gained popularity in theme discovery. These methods independently identify themes from data without the need for pre-labeled datasets. While useful, they also have limitations. Given that topics in topic models are defined as distributions over words, they can be challenging to interpret or may lack coherence, complicating the extraction of meaningful themes (Röder et al. 2015; Lau et al. 2014; Mimno et al. 2011; Chang et al. 2009). Recent advancements have seen the integration of human judgment through a human-in-the-loop approach, which, while enhancing accuracy, suffers from issues of scalability and high resource demand (Pacheco, Islam et al. 2022a, b; Lund et al. 2017; Hu et al. 2011). Recognizing this, our work explores a new approach, integrating the strengths of LLMs (Brown et al. 2020) with a streamlined algorithmic process to efficiently discover and characterize themes in social media messaging. This exploration leads us to the following research questions (RQ), which are crucial for assessing the potential of LLMs to enhance our analysis by providing deeper and more nuanced insights:
@@
To enhance theme discovery, we propose an iterative three-step machine-in-the-loop framework. We define large textual collections as repositories of textual instances (e.g., ads, tweets, posts, documents) where each instance is potentially associated with a set of annotated themes (seed sets from previous studies by Islam et al. (2023); Islam and Goldwasser (2022a)). The seed sets of themes used in this paper are shown in Table 1. A high-level overview of our method is illustrated in Figure 1.
@@
To showcase our approach, we examine two distinct case studies: the climate campaigns and the COVID-19 vaccine campaigns in the United States. In each case, the qualitative researchers apply different theories to ground theme discovery, each associated with a different set of concepts. For climate campaigns, the identification of themes is based on the energy industry and climate change-related stances (Islam et al. 2023). For the COVID-19 vaccine campaigns (Islam and Goldwasser 2022a), the theme discovery is grounded using Moral Foundation Theory (MFT) (Haidt and Graham 2007; Haidt and Joseph 2004). MFT abstracts over specific contentious topics to provide a general framework for understanding human morality. It suggests that there are at least six basic foundations—each with a positive and negative polarity—that explain the similarities and recurrent themes in morality across cultures. These case studies were chosen for their significance to the computational social science (CSS) community, highlighting the challenges of targeted messaging in the digital era. From a machine learning perspective, they present unique challenges due to the specifics of the data, context, and themes involved. Our contributions are summarized as follows:
@@
    We conduct quantitative and qualitative analysis on real-world datasets to demonstrate the effectiveness of our proposed method. We show that our model outperforms the baselines.
@@
    Based on the uncovered themes, we analyze demographic targeting on social media messaging.
@@
Code and the data are available here***https://github.com/tunazislam/latent-themes-llms.
Climate Economy_pro, ClimateSolution, Pragmatism, Patriotism, AgainstClimatePolicy, Economy_clean, FutureGeneration, Environmental, HumanHealth, Animals, SupportClimatePolicy, AltEnergy, PoliticalAffiliation.
COVID-19 GovDistrust, GovTrust, VaccineRollout, VaccineSymptom, VaccineEquity, VaccineStatus, EncourageVaccination, VaccineMandate, VaccineReligion, VaccineEfficacy, VaccineDevelopment, CovidPlan, Vote, VaccineMisinformation, NaturalImmunity.
Table 1: Pre-existing themes from previous studies: Climate campaigns (Islam et al. 2023), COVID-19 vaccine campaigns (Islam and Goldwasser 2022a).
@@
Several studies have been conducted to understand messaging on social media (Islam and et. al 2023; Islam et al. 2023; Islam and Goldwasser 2022b; Capozzi et al. 2021; Silva and Benevenuto 2021; Silva et al. 2020; Serrano et al. 2020; Mejova and Kalimeri 2020; Capozzi et al. 2020; Andreou et al. 2019). Most of the approaches to understanding messaging have relied on broader categories. However, in this paper, we focus on gaining a deeper understanding of messaging.
Previous works (Islam and et. al 2023; Islam et al. 2023; Islam and Goldwasser 2022b; Capozzi et al. 2021) used a predefined set of labels, themes, and arguments to analyze messaging. These were fixed and established based on existing topics or theoretical frameworks. However, previous works often fall short of capturing the nuances of messaging choices. For instance, within the Economy topic, it may neglect to recognize conflicting arguments, such as Argument 1: “Increasing the minimum wage boosts economic growth,” versus Argument 2: “Increasing the minimum wage harms small businesses and reduces employment opportunities.” Recently, Pacheco, Islam et al. (2022a, b) suggested a human-in-the-loop approach and Pacheco, Islam et al. (2023) developed an interactive clustering approach to uncover new themes. In contrast to their work, we use LLMs inference to shape the emergent clusters.
Recently, large language models (LLMs) have made significant progress in learning from prompts via in-context learning (ICL) (Chowdhery et al. 2023; Kojima et al. 2022; Le Scao et al. 2022; Brown et al. 2020). Several studies indicate that LLMs exhibit better performance in tasks traditionally completed by humans (Gilardi, Alizadeh, and Kubli 2023; De Paoli 2023; Dai et al. 2023; Chiang and Lee 2023; Ziems et al. 2024), highlighting a potential to leverage LLMs effectively in our task. Researchers in the fields of qualitative research (QR) and Natural Language Processing (NLP) are currently investigating how LLMs can be utilized for TA. Dai et al. (2023) proposed LLM-in-the-loop approach for TA. Xiao et al. (2023) combined GPT-3 with expert-drafted codebooks for supporting qualitative analysis. Gao and Guo (2024) suggested a collaborative interface for qualitative analysis in their work, utilizing LLMs for generating initial codes and aiding in decision-making processes. In contrast, our framework employs LLMs-in-the-loop to uncover the latent themes in messaging. In this paper, we develop a new set of themes (or codes) specifically tailored for analyzing messaging in a particular domain, which helps speed up the development of domain-specific labels.
Machine-in-the-Loop Framework
@@
Figure 2: Example of an incoherent cluster in Climate.
Figure 3: Example of a merged cluster in Climate.
@@
Figure 4: Prompt template for coherency check (shown as zero-shot).
Figure 5: Prompt template for generating summary (shown as zero-shot).
Figure 6: Prompt template for LLMs Mapper (shown as four-shot for climate).
@@
Naming Cluster
Finally, we prompt LLMs in a zero-shot manner to generate the short descriptive labels/titles for the newly discovered clusters.
Case Method Num. Num. Covered Ads
Study Themes thr $<0.6$ thr $<0.5$ thr $<0.4$ thr $<0.3$
Climate Pre-existing 13 14652 9725 3731 558
+After Iter1 20 18702 14583 8646 2944
+After Iter2 25 18988 15052 9079 3180
COVID-19 Pre-existing 15 7889 6426 3480 771
+After Iter1 20 8852 7627 4737 1302
+After Iter2 23 9092 7898 5038 1590
Table 2: Coverage. THR: Threshold, NUM: Number.
(a) Assignment with LLMs mapper (Climate).
(b) Assignment with SBERT (Climate).
(c) Assignment with LLMs mapper (COVID-19).
(d) Assignment with SBERT (COVID-19).
Figure 7: Confusion matrices after $2^{nd}$ round of iteration + assignment for case studies climate and COVID-19.
@@
The second phase introduces an optional but crucial step – human validation. Two NLP and CSS researchers (age range $30$-$45$, $1$ male, and $1$ female) did one $2$-hour session to assess the clusters formed by the algorithms. The annotators included advanced graduate students and faculty. Their expertise is particularly invaluable in evaluating the nuances of language and context that automated systems might overlook. They scrutinize the merging decisions, the descriptiveness of the clusters, and overall coherence, bridging the gap between computational efficiency and the nuanced understanding of human judgment.
@@
In the final phase, we provide summaries of all clusters (including pre-existing + newly discovered) and prompt LLMs in a few-shot setting to assess whether a new instance aligns with the summarized clusters. To alleviate the issues of being hallucinated, we force LLMs to answer the questions by selecting from the given summarized clusters. The LLMs, equipped with the context from the initial phases, effectively categorize new texts into the relevant clusters (answering RQ2). This phase is crucial for maintaining the dynamism of the clusters, allowing the system to adapt to new data and emerging themes.
Case Studies
We investigate two case studies centered around discussions on social media: (1) climate campaigns and (2) COVID-19 vaccine campaigns. For climate campaigns, we work on the corpus of $21372$ ads and $13$ themes (Table 1) released by Islam et al. (2023). This dataset includes climate-related English ads on Facebook from the US, spanning from January $2021$ to January $2022$. For COVID-19 vaccine campaigns, we use a corpus of $9920$ ads and $15$ themes (Table 1) released by Islam and Goldwasser (2022a) focusing on COVID-19 vaccine-related English ads on Facebook based on US from December $2020$ to January $2022$. Each ad in both corpora includes the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, and distribution of impressions broken down by gender (male, female, unknown), age (seven groups), and location down to the state level in the USA. Additional details about the datasets can be found in the original publications.
Results
Initially, we assign pre-existing themes to the ads using SBERT. Pre-existing themes cover only a small portion of ads, as shown in the first row of Table 2 for climate campaigns and the fourth row of Table 2 for COVID-19 vaccine campaigns. We use a threshold determined by the distance from the cluster centroid to an ad, with shorter distances being preferable. For a threshold of less than $0.5$, there are $11,647$ ads from the climate campaigns and $3,494$ ads from the COVID-19 vaccine campaigns remaining unassigned, which we consider as noisy sets. We then apply k-means clustering to those noisy sets, with $k=29$ for climate and $k=18$ for COVID-19. Following the candidate generation phase, we identify $7$ new themes for climate and $5$ new themes for COVID-19 vaccine campaigns after $1^{st}$ iteration. We use GPT-4 (Achiam et al. 2023) for cluster coherency check, generating cluster summaries, and naming the clusters. In the second iteration, we repeat this process with $6,789$ climate ads and $2,293$ COVID-19 vaccine ads that remained unassigned (noisy sets) from the first iteration (again applying a threshold of $<0.5$). This time, for k-means clustering on noisy sets, we choose $k=20$ for climate and $k=16$ for COVID-19. Iteration $2$ results in the discovery of $5$ additional themes for the climate campaigns and $3$ new themes for the COVID-19 vaccine campaigns. The complete set of final themes from both case studies is presented in Table 3 for each iteration. Summaries of the themes (clusters) for climate campaigns and COVID-19 vaccine campaigns are provided in the Appendix.
Prompt Templates
This section shows the prompt templates used to check cluster coherency (Fig. 4), generate cluster summary (Fig. 5), and map ads to themes (Fig. 6). An example for prompting GPT-4 using three-shot for the COVID-19 dataset is provided herehttps://platform.openai.com/playground/p/AVhxXcExPvQDWP16MT5SdV6u?mode=chat.
Coverage
Our machine-in-the-loop framework successfully discovers new themes that cover a large portion of the messaging for both climate and COVID-19 vaccine campaigns, as detailed in Table 2. In the climate campaigns case study, with a threshold $<0.4$, pre-existing themes initially cover only $17.5\%$ ads. After $1^{st}$ iteration of our machine-in-the-loop approach, we can cover $40.5\%$ (see the second row of Table 2). By the end of the $2^{nd}$ iteration, we manage to cover $42.5\%$ climate ads (refer to the third row of Table 2). For the COVID-19 vaccine campaigns, with the same threshold, pre-existing themes can cover only $35.08\%$ ads. After the first iteration, we are able to cover $47.75\%$ (shown in the fifth row of Table 2) and after $2^{nd}$ iteration, we further cover $50.79\%$ COVID-19 vaccine related ads (see the sixth row of Table 2).
Case Iter. Themes
Study
Climate 1 BidenGasPriceIncrease, AgainstCorporateInterests, GasTax, Deforestation, Carbon, CustomerBasedAltEnergy, FoodSecurity.
2 EnergyAffordabilityandSustainabilityLegislation, EcofriendlyConsumerChoices, PlasticWasteandEnvironmentalImpact, PromoteSustainableTransportation, WaterManagementandSustainability.
COVID-19 1 VaccineRefusalNews, MaskMandate, VaccineBrewIncentives, CommunityServiceByCandidate, AdvocatingUnifiedLiberties.
2 CovidEconomy, UrgentPoliticalAdvocacy, AgainstSocialistPolicy.
Table 3: Resulting themes after each iteration for climate and COVID-19 vaccine campaigns. Iter: Iteration.
@@ Results
(c) Baseline: 15 BERTopic Topics.
(d) Ours: 25 themes after $2^{nd}$ round of iteration.
Figure 8: Correlations between themes and stances for climate case study.
@@
To evaluate the mapping quality of our model, we compare the performance of LLMs mapper with SBERT assignment (Baseline) for ads-to-themes mapping. We select ads that fall under threshold $<0.4$ and then randomly sample $1072$ ads for climate campaigns. For COVID-19 vaccine campaigns, we randomly sample $565$ ads and we do not use a distance threshold in this case study. This approach is deliberately chosen to demonstrate the robustness of our methodology. Typically, ads that are closest to the cluster centroid in the embedding space exhibit higher accuracy and F1 scores for the SBERT assignment. For LLMs mapper assignment, we prompt LLMs in four-shot settings for climate campaigns, and for COVID-19, we use a three-shot prompting to assign ads to themes. However, our LLMs mapper assignment outperforms the SBERT assignment in both case studies, regardless of the distance selection.
Case Study Method Acc. (%) F1 (%)
Climate SBERT Assign. 84.05 79.32
LLMs Mapper 88.15 89.24
COVID-19 SBERT Assign. 41.42 44.83
LLMs Mapper 85.49 81.74
Table 4: Mapping Quality w.r.t Human Judgements. Assign: Assignment, Acc: Accuracy, F1: Macro Average F1 score.
(a) Baseline: 10 LDA Topics.
(b) Baseline: 25 LDA Topics.
(c) Baseline: 15 BERTopic Topics.
(d) Ours: 23 themes after $2^{nd}$ round of iteration.
Figure 9: Correlations between themes and moral foundations for COVID-19 case study.
We ask the annotators to provide a score of $1$ if the ad aligns with the theme, 00 for no alignment. There is no clear definition of alignment, and we let the annotators make this judgment. Results are outlined in Table 4. The resulting mapping from ads-to-themes is fairly accurate (Fig. 7(a)) with respect to human judgments with an accuracy of $88.15\%$ and macro average F1 score of $89.24\%$ for the climate case study. On the other hand, we observe a noticeable improvement (Fig. 7(c)) in accuracy ($85.49\%$) and F1 score ($81.74\%$) for COVID-19 case study. Fig. 7 shows the confusion matrix after $2^{nd}$ round of iteration of LLMs mapper and SBERT assignment. For climate, Fig. 7(b) illustrates the SBERT mapper assignments, which are comparable to those of the LLMs mapper (Fig. 7(a)) due to ad selection from closest to the cluster centroid (threshold $<0.4$). However, a significant difference is observed between the LLMs mapper (Fig. 7(c)) and the SBERT mapper (Fig. 7(d)) in the COVID-19 case study, attributed to the random selection of ads independent of the distance threshold.
Case Study Age Group State Theme Entity
Climate 13-24 CA Environmental Young People, Climate Change Denier, Climate Movement.
TX AltEnergy Natural Gas Systems, Coal Energy Systems, Nuclear Energy Systems, Wind Turbines, Solar Panels.
25-54 CA Carbon Carbon Emissions, Sacramento Municipal Utility District.
TX AltEnergy Climate Change, Renewable Energy Plan.
55+ CA Animals Western Burrowing Owls, Aramis Industrial Solar Power Plant.
TX SupporClimPolicy U.S. Senators or Congressman, Citizens’ Climate Lobby.
COVID 13-24 TX VaccineRollout Kids/Children, Team/Teammates.
-19 25-54 FL UrgPoliticalAdvc Ron DeSantis, Dr. Joseph Ladapo, Surgeon General, Floridians.
55+ TX EncourageVax Seniors, Pfizer, Elders.
FL VaccineRollout Governor Ron DeSantis, Seniors, Bosters, Johnson & Johnson vaccine.
Table 5: Most mentioned entities, claim, and theme of targeted ads for young, working-age, older population of CA and TX from climate dataset. TX and FL from COVID-19 dataset. SupClimPolicy: SupportClimatePolicy, UrgPoliticalAdvc: UrgentPoliticalAdvocacy, EncourageVax: EncourageVaccination.